Let X be a random variable that only takes the values 0,1,2, 3, 4, .. Prove that x3 n=1
Answer #1

E(X) = ∑ n⋅P(X=n)

= P(X=1) + P(X=2)+P(X=2) + P(X=3)+P(X=3)+P(X=3) + P(X=4)+P(X=4)+P(X=4)+P(X=4) + ...

= p1+2p2+3p3+4p4+5p5+6p6+⋯

=Pr(X>0)+p2+2p3+3p4+4p5

=Pr(X>0)+Pr(X>1)+p3+2p4+3p5+⋯

=Pr(X>0)+Pr(X=1)+Pr(X=2)+p4+2p5+⋯

= ∑P(X>=n)

Hence, proved